Assume that the required rate of return on a stock is 15.6% and the expected return on the market is 14%. What is the risk-free rate if the stock has a beta of 1.2? *
1
Expert's answer
2020-06-10T18:56:58-0400

E(ri ) = rf + β1 [ E (rm ) - rf )


15.6 = rf + 1.2 (0.14 - rf )


0.2 = rf + 0.168 - 1.2 rf


0.2 rf = -0.032


rf = -0.16


rf =